# SoK: LLM-based Log Parsing

### v1

@@ 4.8. LLM Usage (PS-4)
  • Target log(s): Contains the log(s) to be parsed. The approaches (Vaarandi and Bahsi, 2024; Xiao et al., 2024; Huang et al., 2024; Xu and Gau, 2024; Ma et al., 2024b; Pei et al., 2024) provide multiple target logs at once in the query, while the rest provide single logs. Providing multiple logs at once can help the LLM in abstracting variable and static parts from the logs (Pei et al., 2024).
Le et al. (Le and Zhang, 2023a) find that with a simple prompt, where no context is provided nor a detailed instruction is given, GPT-3.5 is hardly able to understand the concept of log parsing. Sun et al. (Sun et al., 2024a) achieve a significantly higher PA ($20\%$ to $60\%$) with a prompt that contains extraction rules for direct parsing (on HDFS and BGL with GPT-3.5 and GPT-4) than without extraction rules.
@@ 4.9. Template Revision (PS-5)
In total, $14$ papers revise templates in a post-processing step. All works that correct their templates in a postprocessing step also use a cache (PS-3). The works (Zhang et al., 2025; Zhong et al., 2024; Wu et al., 2024; Liu et al., 2024) employ the LLM to revise templates per prompt: Lemur (Zhang et al., 2025) uses CoT in three dialogue rounds for revising semantically similar templates for potential merge. First, the structure is revised, then the semantics, and in a final round the LLM is asked for a solution based on the first two rounds. LogParser-LLM (Zhong et al., 2024) and LogBabylon (Karanjai et al., 2024) generate a parsing tree from the templates during parsing. In case, a loose match is identified the LLM decides whether the template can be merged or not, leading to an update of the template tree, or the creation of a new tree root node. The works (Liu et al., 2024; Ma et al., 2024b; Wu et al., 2024) re-prompt the logs to the LLM, with the same prompt they were initially parsed, to the LLM if they do not match all the related logs until all logs are matched (Liu et al., 2024) or if a certain number of re-prompts is exceeded (Ma et al., 2024b). AdaParser (Wu et al., 2024) also revises templates if their variables contain certain keywords linked to exceptions, failures or interrupts, which they state should not be abstracted with a wildcard (Wu et al., 2024).
Updating the previously parsed templates can adapt the parser to changes in the computer system and help correct faulty templates. Templates are typically faulty because they have static parts of the log identified as a variable, or variable parts were identified as static. Faulty templates can be identified by high similarity with an existing template (Jiang et al., 2024a; Zhang et al., 2024a; Zhong et al., 2024), if not all logs within a similarity cluster can be matched (Xiao et al., 2024), or if the centroids of log clusters move closer together through newer incoming logs added to clusters (Xu and Gau, 2024). Templates can be merged with templates that partially match with an existing template, by traversing a prefix tree (Jiang et al., 2024a; Karanjai et al., 2024; Zhong et al., 2024), LCS (Zhang et al., 2024a; Zhi et al., 2024), implicitly by reentering the logs into the parsing queue (Xiao et al., 2024), or by monitoring whether multiple paths of a prefix tree join together in a subsequent node. A previously parsed template can also be replaced with a new and more permissive template (Zhi et al., 2024; Vaarandi and Bahsi, 2024). More permissive in this context means that more variables were identified for the newer template.
@@ 4.11. Metrics (R-2)
Since the correctness evaluation of templates is not straightforward, there is a strong variety of different evaluation metrics. The set of all used metrics is $\{ED,FGA,FTA,GA,PA,PTA,RTA,\texttt{other}\}$. Except for “other”, these are the traditional metrics that have been widely used (Jiang et al., 2024b; Khan et al., 2022). If we compute the Jaccard Index of the used metrics and all available metrics and average the results we get roughly $0.31$, which means that on average only $31\%$ of the full range of metrics is used (including “other” as a metric). More specifically, $5$ publications use only PA, or PA and GA, $7$ use ED, $9$ publications use at least one metric that is insensitive to class imbalance (FGA, FTA), $8$ publications use metrics other than the traditional ones.
The metrics cover different characteristics of correctness. It is therefore important for evaluating parsers to cover the relevant evaluation aspects with these metrics. For instance, Jiang et al. (Jiang et al., 2024b) proposed using FGA and FTA since they are insensitive to class imbalance, while Khan et al. (Khan et al., 2022) recommend using GA, PA, RTA and PTA to cover all aspects (FTA is the harmonic mean of RTA and PTA). For example, the anomaly detection tool, DeepLog (Du et al., 2017), detects anomalies solely based on the event ID. Each event ID corresponds to a unique template. Consequently, it is indifferent if the parsed templates are correct at the template level as long as they are grouped together correctly, which is indicated by a high score for GA. Since detection algorithms can focus on a variety of data characteristics (Beck et al., 2024), it is important to report metrics that cover these characteristics.
@@ 4.12. Model (R-3)
Table 3. Overview of used LLMs.
@@ 4.13. Code (R-4)
From our selection, $16$ papers make their implementation available and provide links to their code repositories. At the time of writing, two of them (Huang et al., 2024; Sun et al., 2024b) lead to an empty or non-existing repository, but consider that both are preprint versions. The rest did not make their code open-source. A study by Olzewski et al. (Olszewski et al., 2023) analyzed the reproducibility and replicability of $750$ machine learning papers and their codebases and datasets from Tier 1 security conferences between the years 2013 and 2022. They found that about $59\%$ of papers did not provide any reproducible artifacts. In our study, $45\%$ of papers do not provide any reproducible artifacts.
@@ 4.13.1. Code Quality
A large proportion of the approaches do not provide code and therefore can not be used for the evaluation in Sec. 6. Even with available code, some do not provide the comprehensive functionality to replicate the processes described in their papers and comprehensively correcting the code of others is out of scope for this work. At the time of review (February 14th, 2025) the code of Lemur [^4] https://github.com/zwpride/lemur; accessed 14-February-2025. (Zhang et al., 2025) lacks a script for selecting the previously parsed templates for the subsequent CoT template merging process. The code of LogGenius [^5] https://github.com/huashengyihao/LogGenius; accessed 14-February-2025. (Yu et al., 2024) runs into multiple errors due to missing folders and files, while there is no README file providing instructions. The LogEval repository [^6] https://github.com/LinDuoming/LogEval; accessed 14-Febrary-2025. (Cui et al., 2024) does not provide the scripts for parsing nor instructions. LogPrompt [^7] https://github.com/lunyiliu/LogPrompt; accessed 14-February-2025. (Liu et al., 2024) encounters multiple errors due to simple typos. Furthermore, they describe three different prompt strategies, but it is not clear which they used for their result nor do they explain it in the instruction file. Moreover, they do not provide the functionality to run the parsing process except for the self-prompt setting. SelfLog (Pei et al., 2024) proposes a self-evolutionary approach, in which they “retrieve the most similar historical logs and their templates from the data through an Approximate Nearest Neighbors (ANN) search, serving as the corpus for In-Context Learning (ICL)(Pei et al., 2024). One could think that this means they update the database incrementally after each newly parsed log. However, in the code [^8] https://github.com/CSTCloudOps/SelfLog; accessed 14-February-2025. the script provided for the effectiveness evaluation does not contain the self-evolution functionality. They provide a second file for online parsing that does provide this functionality but it does not work ad hoc due to missing files and missing function parameters. Also, by default they do not use ANN but cosine similarity for retrieval which is somewhere hidden in the code (not an adjustable parameter). The code repository of SuperLog [^9] https://github.com/J-York/SuperLog; accessed 14-Febrary-2025 (Ji et al., 2024) only provides code for LLM training, but not the described framework for parsing. Moreover, we found that many repositories provide sparse instructions, which further impedes reproducibility and application, especially when the execution scripts do not work ad hoc.
@@ 4.13.2. Licenses
Table 4. Licenses of the approaches. Most approaches do not provide a license. Please note, that especially for preprint papers the license might change or, if absent, be added in the future (status 7-March-2025).
@@
4.14. Notable Mentions and Promising Techniques
Besides the features we defined in Sec 3.2, we identified other promising techniques and ideas, worth mentioning in the reviewed papers. For instance, some approaches specifically focus on highlighting the commonalities and variabilities of logs. Lemur (Zhang et al., 2025), LUNAR (Huang et al., 2024), OpenLogParser (Ma et al., 2024b), and LogBatcher (Xiao et al., 2024) cluster the logs into multilevel clusters for parsing while the supervised parsers DivLog (Xu et al., 2024) and LILAC (Jiang et al., 2024a) use this concept for sampling labeled logs. The coarse-grained clusters thereby capture the commonality of the logs, that, in the best case, all belong to the same log group (i.e. the same event), and the fine-grained clusters should highlight the variabilites within that group. This idea is also used by LogShrink (Li et al., 2024) for compressing large log files.
Xiao et al. (LogBatcher) (Xiao et al., 2024) report to use the heuristic rules proposed by Khan et al. (Khan et al., 2022) to correct templates. This includes measures, such as combining subsequent wildcards, like <*><*>, into a single one <*>. This technique is also used by(Xu and Gau, 2024; Jiang et al., 2024a; Xu et al., 2024) and other work outside of our selection (Liu et al., 2022; Le and Zhang, 2023b). By manually inspecting some the templates generated by the parsers used for the evaluation, we found that a significant proportion of templates would have been considered correct by the evaluation functions, if some of these steps would have been applied.
@@ 5. Evaluation Setup
In terms of evaluation, the existing literature regarding LLM-based log parsing misses a common ground. While there are some commonalities regarding single features like used datasets, used evaluation metrics, or used models, the combination of these is highly varying between the approaches - see feature R-1, R-2, and R-3 in Table 2. As a consequence, we create a benchmark featuring a subset of $7$ out of the $29$ approaches. To counteract randomness from sampling, LLM output and runtime fluctuations, we run each parser $3$ times take the average.
@@ 5.1. Selection for the Benchmark
For the benchmark we only select parsers which made their code publicly available — see feature R-4 in Table 2. We distinguish between approaches that solely focus on methods changing model parameters, like fine-tuning and pre-training, and approaches that apply the LLM as a tool within a framework. Prior studies showed that fine-tuned (Mehrabi et al., 2024) or pre-trained (Ji et al., 2024) models can achieve SotA accuracy scores without complex pre- and post-processing modules for RAG (PS-2), caching (PS-3), and template revision (PS-5). However, it is clear that improved quality of the LLM’s output leads to overall improved output quality of the whole framework. It is therefore meaningful to separate model-centric and model-wrapping approaches. For our evaluation, we focus on the latter and therefore do not include LLMParser (Ma et al., 2024a), SuperLog (Ji et al., 2024) and the work of Mehrabi et al. (Mehrabi et al., 2024) since they do not focus on frameworks but rather on the LLM itself, even tough they provide their code or model, respectively. We also exclude Semirald (Sun et al., 2024a) from the evaluation since its parsing approach consists only of naive queries to an LLM without other processing steps or model training.
In Sec. 4.13.1 we described that some approaches provide code, but do not provide functioning parsers. They are consequently not included in the benchmark (except for LogPrompt).
Given the above-stated matters the remaining approaches featured in the benchmark are LILAC (Jiang et al., 2024a), OpenLogParser (Ma et al., 2024b), LogBatcher (Xiao et al., 2024), SelfLog (Pei et al., 2024), DivLog (Xu et al., 2024), LLM-TD (Vaarandi and Bahsi, 2024), LogPrompt (Liu et al., 2024). To a large extent, this selection represents the state of the art of LLM-based log parsing frameworks, as their approaches cover most of the aspects of this field. Unfortunately, due to the aforementioned selection criteria, this benchmark could not include any work using LLMs as pre- or post-processors.
@@ 5.2. Datasets and Baseline
Following previous work (Xiao et al., 2024; Astekin et al., 2024b, a; Zhi et al., 2024; Ma et al., 2024a; Le and Zhang, 2023a, b) we take the corrected version of LogHub from Khan et al. (Khan et al., 2022) as the default dataset for our evaluation. It consists of $16$ datasets with ground-truth templates from different domains such as distributed and supercomputer systems, and server application. Note that we found minor errors in the “Content” column of the corrected version, mostly additional character spaces, that led to errors when evaluating. We exchanged this column (during runtime) with the “Content” column from the original LogHub version (Zhu et al., 2023) (only the templates are different but the content is the same). In Sec. 6.1, we also use the original LogHub version (Zhu et al., 2023) which contains the non-corrected templates. For the runtime evaluation in Sec. 6.2 we use the large-scale datasets from LogHub-2.0 (Jiang et al., 2024b).
@@
Since the LogHub datasets are so widely used for the evaluation of log parsers (see Sec. 4.5) there is a tendency that the parsers work especially well on these datasets but not on others. To simulate a use case closer to a real life example, we extend the $16$ datasets by a custom dataset based on AIT Log Dataset V2 (Landauer et al., 2023b, 2021). The dataset contains $2000$ audit logs (from russellmitchell/gather/intranet_server/logs/audit) with $8$ unique templates. The logs were manually annotated by ourselves, adhering to the style of the templates from the corrected LogHub version (Khan et al., 2022). Audit logs track system activities, user actions, and changes for security, compliance, and troubleshooting. The corresponding files can be found in our code repository.
@@
For the baseline of the evaluation, we select five non-LLM-based parsers from the LogPAI logparser repository [^10] https://github.com/logpai/logparser; accessed 21-March-2025., namely AEL (Jiang et al., 2008), SPELL (Du and Li, 2016), Drain (He et al., 2017), ULP (Sedki et al., 2022) and Brain (Yu et al., 2023). AEL, SPELL and Drain have been widely used in the log parsing research and their performance has been thoroughly investigated by (Jiang et al., 2024b; Zhang et al., 2023; Khan et al., 2022). Their papers were published in 2008 (AEL), 2016 (SPELL) and 2017 (Drain). ULP and Brain are newer advancements from the years 2022 and 2023. For the baseline parsers the default configuration was used, but consider that the default configuration of all these parsers contains the log format, regular expressions (but empty for some datasets), and two or more other parameters that are specific to each of the LogHub datasets. For the Audit dataset (Landauer et al., 2023b, 2021) the log format “type=<Type> msg=audit(<Time>): <Content>” was used and the regular expression parameter was left empty. The other parameters were simply the parameter configuration that is used most often for the other datasets. The parameters of the configuration were not specifically tuned.

### v2

@@ 5.8. LLM Usage (PS-4)
  • Target log(s): Contains the log(s) to be parsed. The approaches (Vaarandi and Bahsi, 2024; Xiao et al., 2024; Huang et al., 2024; Xu and Gau, 2024; Ma et al., 2024b; Pei et al., 2024) provide multiple target logs at once in the query, while the rest provide single logs. Providing multiple logs at once can help the LLM in abstracting variable and static parts from the logs (Pei et al., 2024; Xiao et al., 2024).
Le et al. (Le and Zhang, 2023a) find that with a simple prompt, where no context is provided nor a detailed instruction is given, GPT-3.5 is hardly able to understand the concept of log parsing. Sun et al. (Sun et al., 2024a) achieve a significantly higher PA ($20\%$ to $60\%$) with a prompt that contains extraction rules for direct parsing (on the HDFS and BGL datasets from LogHub (Zhu et al., 2023) with GPT-3.5 and GPT-4) than without extraction rules.
@@ 5.9. Template Revision (PS-5)
In total, $14$ papers revise templates in a post-processing step. All works that correct their templates in a postprocessing step also use a cache (PS-3). The works (Zhang et al., 2025; Zhong et al., 2024; Wu et al., 2024; Liu et al., 2024) employ the LLM to revise templates per prompt: Lemur (Zhang et al., 2025) uses CoT in three dialogue rounds for revising semantically similar templates for potential merge. First, the structure is revised, then the semantics, and in a final round the LLM is asked for a solution based on the first two rounds. LogParser-LLM (Zhong et al., 2024) and LogBabylon (Karanjai et al., 2024) generate a parsing tree from the templates during parsing. In case, a loose match is identified the LLM decides if the template should be merged, leading to an update of the template tree, or the creation of a new tree root node. The works (Liu et al., 2024; Ma et al., 2024b; Wu et al., 2024) re-prompt the logs to the LLM, with the same prompt they were initially parsed, to the LLM if they do not match all the related logs until all logs are matched (Liu et al., 2024) or if a certain number of re-prompts is exceeded (Ma et al., 2024b). AdaParser (Wu et al., 2024) also revises templates if their variables contain certain keywords linked to exceptions, failures or interrupts, which they state should not be abstracted with a wildcard (Wu et al., 2024).
Updating the previously parsed templates can adapt the parser to changes in the computer system and help correct faulty templates. Templates are typically faulty because they have static parts of the log identified as a variable, or variable parts were identified as static. Faulty templates can be identified by high similarity with an existing template (Jiang et al., 2024a; Zhang et al., 2024a; Zhong et al., 2024), if not all logs within a similarity cluster can be matched (Xiao et al., 2024), or if the centroids of log clusters move closer together through newer incoming logs added to clusters (Xu and Gau, 2024). Templates can be merged with templates that partially match with an existing template, by traversing a prefix tree (Jiang et al., 2024a; Karanjai et al., 2024; Zhong et al., 2024), LCS (Zhang et al., 2024a; Zhi et al., 2024), implicitly by reentering the logs into the parsing queue (Xiao et al., 2024), or by monitoring whether multiple paths (templates) of a prefix tree join together in a subsequent node (Pei et al., 2024). A previously parsed template can also be replaced with a new and more permissive template (Zhi et al., 2024; Vaarandi and Bahsi, 2024). More permissive in this context means that more variables were identified for the newer template.
@@ 5.10. Datasets (R-1)
Using custom datasets alongside LogHub datasets offers a broader view of parser performance but can affect result comparability. The studies (Khan et al., 2022; Fu et al., 2022; Zhang et al., 2024a) show that parsers often perform well on LogHub datasets (Zhu et al., 2023) or its descendants (Jiang et al., 2008; Khan et al., 2022) but poorly on custom datasets. Additionally, it is not clear whether the open-source log datasets used are known to the LLMs from their extensive pretraining phases. However, the ablative studies of (Jiang et al., 2024a; Xu et al., 2024; Xiao et al., 2024) report significant performance improvements with ICL compared to direct zero-shot parsing, arguing that this implies a low likelihood of memorization of single log templates for GPT-3 and GPT-3.5-Turbo.
@@ 5.11. Metrics (R-2)
Since the correctness evaluation of templates is not straightforward, there is a strong variety of different evaluation metrics. The set of all used metrics is $\{ED,FGA,FTA,GA,PA,PTA,RTA,\texttt{other}\}$. Except for “other”, these are the traditional metrics that have been widely used (Jiang et al., 2024b; Khan et al., 2022). If we compute the Jaccard Index of the used metrics and all available metrics and average the results we get roughly $0.33$, which means that on average only $33\%$ of the full range of metrics is used (including “other” as a metric). More specifically, $5$ publications use only PA, or PA and GA, $6$ use ED, $12$ publications use at least one metric that is insensitive to class imbalance (FGA, FTA). Also, $5$ publications use metrics other than traditional ones. For example, the creators of LogBatcher (Xiao et al., 2024) report the normalized Edit Distance (NED) (Marzal and Vidal, 1993) which computes the mean ED of all template pairs compared in the data set. In (Astekin et al., 2024b) Astekin et al. report the number of unique templates as a means of LLM determinism, while in (Astekin et al., 2024a) they report variations of metics based on ED and LCS. The creators of LogParser-LLM (Zhong et al., 2024) propose metrics based on Granularity Distance, the minimum number of operations necessary to transform one parsing result into the other. An operation is thereby defined as converting a static part of the parsed template to a variable or vice versa.
The metrics cover different characteristics of correctness. It is therefore important for evaluating parsers to cover the relevant evaluation aspects with these metrics. For instance, Jiang et al. (Jiang et al., 2024b) proposed using FGA and FTA since they are insensitive to class imbalance, while Khan et al. (Khan et al., 2022) recommend using GA, PA, RTA and PTA to cover all aspects (FTA is the harmonic mean of RTA and PTA). For example, the anomaly detection tool, DeepLog (Du et al., 2017), detects anomalies solely based on the event ID. Each event ID corresponds to a unique template. Consequently, it is indifferent if the parsed templates are correct at the template level as long as they are grouped together correctly, which is indicated by a high score for GA. Since downstream tasks, like anomaly detection, can focus on a variety of data characteristics (Beck et al., 2024), it is important to report metrics that cover these characteristics.
Previous work (Petrescu et al., 2023; Liu et al., 2022; Nedelkoski et al., 2021) found that the state-of-the-art log parsers are often evaluated with insufficient selections of evaluation metrics. While their work concerns the conventional non-LLM based approaches, we find that many LLM-based approaches carry on with this manner. More specifically, we consider using only PA or the combination of the grouping metrics GA and FGA as insufficient. The highly varying scores between the metrics of different parsers in (Jiang et al., 2024b; Khan et al., 2022) but also in our own evaluation in Sec. 7 show how important a versatile evaluation metric selection is. Focusing only on a single aspect misleads estimations of a parser’s performance.
@@ 5.12. Model (R-3)
Table 3. Overview of LLMs used.
@@ 5.13. Code (R-4)
From our selection, $16$ papers make their implementation available and provide links to their code repositories. At the time of writing, two of them (Huang et al., 2024; Sun et al., 2024b) lead to an empty or non-existing repository, but consider that both are preprint versions. The rest did not make their code open-source. A study by Olzewski et al. (Olszewski et al., 2023) analyzed the reproducibility and replicability of $750$ machine learning papers and their codebases and datasets from Tier 1 security conferences between the years 2013 and 2022. They found that about $59\%$ of the papers did not provide any reproducible artifacts. In our study, $45\%$ of the reviewed papers do not provide any reproducible artifacts (including preprints).
@@ 5.13.1. Code Quality
A large proportion of the approaches do not provide code and therefore can not be used for the evaluation in Sec. 7. Even with available code, some do not provide the comprehensive functionality to replicate the processes described in their papers and comprehensively correcting the code of others is out of scope for this work. At the time of review (February 14th, 2025) the code of Lemur [^3] https://github.com/zwpride/lemur; accessed 14-February-2025. (Zhang et al., 2025) lacks a script for selecting the previously parsed templates for the subsequent CoT template merging process. The code of LogGenius [^4] https://github.com/huashengyihao/LogGenius; accessed 14-February-2025. (Yu et al., 2024) runs into multiple errors due to missing folders and files, while there is no README file providing instructions. The LogEval repository [^5] https://github.com/LinDuoming/LogEval; accessed 14-Febrary-2025. (Cui et al., 2024) does not provide the scripts for parsing nor instructions. LogPrompt [^6] https://github.com/lunyiliu/LogPrompt; accessed 14-February-2025. (Liu et al., 2024) encounters multiple errors due to simple typos (e.g. a variable that seems to be undefined but it is actually spelled differently than in its definition). Furthermore, they describe three different prompt strategies, but it is not clear which they used for their result nor do they explain it in the instruction file. Moreover, they do not provide the functionality to run the parsing process except for the self-prompt setting. SelfLog (Pei et al., 2024) proposes a self-evolutionary approach, in which they “retrieve the most similar historical logs and their templates from the data through an Approximate Nearest Neighbors (ANN) search, serving as the corpus for In-Context Learning (ICL)(Pei et al., 2024). One could think that this means they update the database incrementally after each newly parsed log. However, in the code [^7] https://github.com/CSTCloudOps/SelfLog; accessed 14-February-2025. the script provided for the effectiveness evaluation does not contain the self-evolution functionality. They provide a second file for online parsing that does provide this functionality but it does not work ad hoc due to missing files and missing function parameters. Also, by default they do not use ANN but cosine similarity for retrieval which is somewhere hidden in the code (not an adjustable parameter). The code repository of SuperLog [^8] https://github.com/J-York/SuperLog; accessed 14-Febrary-2025 (Ji et al., 2024) only provides code for LLM training, but not the described framework for parsing. Moreover, we found that many repositories provide sparse instructions, which further impedes reproducibility and application, especially when the execution scripts do not work ad hoc.
@@ 5.13.2. Licenses
Table 4. Licenses of the approaches with open-source implementations. Most approaches do not provide a license. Please note, that especially for preprint papers the license might change or, if absent, be added in the future (status 7-March-2025).
@@
5.14. Notable Mentions
Besides the features we defined in Sec 4.2, we identified other promising techniques and ideas, worth mentioning in the reviewed papers. For instance, some approaches specifically focus on highlighting the commonalities and variabilities of logs. Lemur (Zhang et al., 2025), LUNAR (Huang et al., 2024), OpenLogParser (Ma et al., 2024b), and LogBatcher (Xiao et al., 2024) cluster the logs into multilevel clusters for parsing while the supervised parsers DivLog (Xu et al., 2024) and LILAC (Jiang et al., 2024a) use this concept for sampling labeled logs. The coarse-grained clusters thereby capture the commonality of the logs, that, in the best case, all belong to the same log group (i.e. the same event), and the fine-grained clusters should highlight the variabilites within that group. This idea is also used by LogShrink (Li et al., 2024b) for compressing large log files.
Xiao et al. (LogBatcher) (Xiao et al., 2024) report to use the heuristic rules proposed by Khan et al. (Khan et al., 2022) to correct templates. This includes measures, such as combining subsequent wildcards, like <*><*>, into a single one <*>. This technique, or similar, is also used by (Xu and Gau, 2024; Jiang et al., 2024a; Xu et al., 2024) and other work outside of our selection (Liu et al., 2022; Le and Zhang, 2023b). By manually inspecting some the templates generated by the parsers used for the evaluation, we found that a significant proportion of templates would have been considered correct by the evaluation functions, if some of these steps would have been applied.
@@ 6. Evaluation Setup
In terms of evaluation, the existing literature regarding LLM-based log parsing misses a common ground. While there are some commonalities regarding single features, such as used datasets, used evaluation metrics, or used models, the combination of these is highly varying between the approaches - see feature R-1, R-2, and R-3 in Table 2. As a consequence, we create a benchmark comparing a subset of $7$ out of the $29$ approaches. To counteract randomness from sampling, LLM output and runtime fluctuations, we run each parser $3$ times and take the average.
@@ 6.1. Selection for the Benchmark
For the benchmark we only select parsers which made their code publicly available — see feature R-4 in Table 2. We distinguish between approaches that solely focus on methods changing model parameters, like fine-tuning and pretraining, and approaches that apply the LLM as a tool within a framework. Prior studies showed that fine-tuned (Mehrabi et al., 2024) or pretrained (Ji et al., 2024) models can achieve state-of-the-art accuracy scores without pre- and post-processing modules for RAG (PS-2), caching (PS-3), and template revision (PS-5). However, it is clear that improved quality of the LLM’s output leads to overall improved output quality of the whole framework. It is therefore meaningful to separate model-centric and model-wrapping approaches — especially, in the rapidly evolving landscape of LLMs, where more performant models are introduced on a daily basis. For our evaluation, we limit our scope and focus on these frameworks and therefore do not include the fine-tuning or pretraining approaches LLMParser (Ma et al., 2024a), SuperLog (Ji et al., 2024), and the work of Mehrabi et al. (Mehrabi et al., 2024), even tough they provide their code or model. We also exclude Semirald (Sun et al., 2024a) since it does not provide a log parsing framework and focuses rather on anomaly detection. In Sec. 5.13.1, we described that some approaches provide code, but do not provide functioning parsers. They are consequently not included in the benchmark (except for LogPrompt since its typos could be resolved by exchanging single characters).
Given the above-stated matters the remaining approaches featured in the benchmark are LILAC (Jiang et al., 2024a), OpenLogParser (Ma et al., 2024b), LogBatcher (Xiao et al., 2024), SelfLog (Pei et al., 2024), DivLog (Xu et al., 2024), LLM-TD (Vaarandi and Bahsi, 2024), LogPrompt (Liu et al., 2024). To a large extent, this selection represents the state-of-the-art of LLM-based log parsing frameworks, as their approaches cover most of the aspects of this field. Unfortunately, due to the aforementioned selection criteria, this benchmark could not include any work using LLMs as pre- or post-processors.
@@ 6.2. Datasets and Baseline
Following previous work (Xiao et al., 2024; Astekin et al., 2024b, a; Zhi et al., 2024; Ma et al., 2024a; Le and Zhang, 2023a, b) we take the corrected version of LogHub from Khan et al. (Khan et al., 2022) as the default dataset for our evaluation. It consists of $16$ datasets with ground-truth templates from different domains such as distributed and supercomputer systems, and server application. Note that we found minor errors in the “Content” column of the corrected version, mostly additional character spaces, that led to errors when evaluating. We exchanged this column (during runtime) with the “Content” column from the original LogHub version (Zhu et al., 2023) (only the templates are different but the content is the same). In Sec. 7.1, we also use the original LogHub version (Zhu et al., 2023) which contains the non-corrected templates. For the runtime evaluation in Sec. 7.2 we follow previous work (Jiang et al., 2024a) and only report the runtime of the large-scale datasets from LogHub-2.0 (Jiang et al., 2024b). This choice is motivated by the need for a more extensive data as the $2000$ samples from the original LogHub are insufficient for a comprehensive runtime analysis. The results for the runtime on the LogHub datasets are available at (Beck, 2025).
@@
Since the LogHub datasets are so widely used for the evaluation of log parsers (see Sec. 5.5) there is a tendency that the parsers work especially well on these datasets but not on others. To simulate a use case closer to a real life example, we extend the $16$ datasets by a custom dataset based on AIT Log Dataset V2 (Landauer et al., 2023b, 2021). The dataset contains $2000$ audit logs (from russellmitchell/gather/intranet_server/logs/audit) with $8$ unique templates. This dataset will be referred to as Audit in the following. The logs were manually annotated by ourselves, adhering to the style of the templates from the corrected LogHub version (Khan et al., 2022). Audit logs track system activities, user actions, and changes for security, compliance, and troubleshooting. The corresponding files can be found in our code repository (Beck, 2025).
@@
For the baseline of the evaluation, we select five non-LLM-based parsers from the LogPAI logparser repository [^9] https://github.com/logpai/logparser; accessed 21-March-2025., namely AEL (Jiang et al., 2008), SPELL (Du and Li, 2016), Drain (He et al., 2017), ULP (Sedki et al., 2022) and Brain (Yu et al., 2023). AEL, SPELL and Drain have been widely used in the log parsing research and their performance has been thoroughly investigated by (Jiang et al., 2024b; Zhang et al., 2023; Khan et al., 2022). Their papers were published in 2008 (AEL), 2016 (SPELL) and 2017 (Drain). ULP and Brain are newer advancements from the years 2022 and 2023. The baseline parsers use the standard configuration of the literature (Xiao et al., 2024; Jiang et al., 2024a; Zhu et al., 2023; Jiang et al., 2024b). The default configuration of all these parsers contains the log format, regular expressions (but empty for some datasets), and up to two other parameters that are specific to each of the LogHub datasets. For the Audit dataset (Landauer et al., 2023b, 2021) the log format “type=<Type> msg=audit(<Time>): <Content>” was used and the regular expression parameter was left empty. The other parameters for the Audit dataset were chosen through grid search by the best score for GA. We used the full dataset for this hyperparameter tuning to show the best possible performance of the baseline.